OpenStudy (anonymous):

Simplify: 4 x 2x2 + 5 – 2x +2 2x2 + 5 – 2x

OpenStudy (anonymous):

Can anyone pls help me with the attached question. Thanks

OpenStudy (anonymous):

2^x = t \[\frac{t^2*2^5*4 - t*2^2}{t^2*2^5 -t}\]

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

yes i got it! thanks! ^^

OpenStudy (anonymous):

\[4\frac{t^5-t}{t^5-t}\] \[4\]